OpenStudy (anonymous):

what is the derivative of:

OpenStudy (anonymous):

\[A(x)=(2x)(\sqrt{r ^{2}-x ^{2}})\]

OpenStudy (anonymous):

where \(r\) is constant I guess?

OpenStudy (anonymous):

This is a mix of the product rule and the chain rule then.

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Product rule is \[(uv)'= u'v + v'u\]

OpenStudy (anonymous):

For the radical expression you need to apply the chain rule.

OpenStudy (anonymous):

okay..thanks! :)

OpenStudy (anonymous):

welcome, gives it a try, we can afterwards check your answer together or you can enter it in wolframalpha.

OpenStudy (anonymous):

okay..later! :)

OpenStudy (anonymous):

is the answer: \[A'=(2r ^{2}+2xr) \div (\sqrt{r ^{2}-x ^{2}})\] ?

OpenStudy (anonymous):

i'm solving for a calculus problem.. so i must equate that derivative to zero..and solve for x... how?

OpenStudy (anonymous):

i'm confused

OpenStudy (anonymous):

how'd you get that? @Spacelimbus

OpenStudy (anonymous):

let me try that again step by step before i make silly mistakes.

OpenStudy (anonymous):

\[\Large A'= 2\sqrt{r^2-x^2} - \frac{2x^2}{\sqrt{r^2-x^2}} \]

OpenStudy (anonymous):

which is equal to \[ \Large A'= \frac{2(r^2-2x^2)}{\sqrt{r^2-x^2}} \]